The 2008 balance sheet of Maria's Tennis Shop, Inc., showed $2.65 million in long-term debt, $720,000 in the common stock account, and $6.25 million in the additional paid-in surplus account. The 2009 balance sheet showed $3.6 million, $955,000, and $8.35 million in the same three accounts, respectively. The 2009 income statement showed an interest expense of $260,000. The company paid out $670,000 in cash dividends during 2009. If the firm's net capital spending for 2009 was $740,000, and the firm reduced its net working capital investment by $125,000, the firm's 2009 operating cash flow, or OCF? Multiple Choice $-1,740,000 $-2,355,000 $-4,310,000 $2,605,000 $-3,080,000
Answer #1

Cash Flow to  Creditors + Cash Flow Equity holders = Interest - New Debt = 260,000 - (3,600,000-2,650,000) = -690,000

Cash Flow Equity holders = (Dividends - New Equity) = 670,000 - ((955,000+8,350,000)-(625,000+720,000)) = -1,665,000

Cash Flow from Assets = Cash Flow from Creditors + Cash Flow Equity holders = -690,000 -1,665,000 = -2,355,000

CFA =OCF - Increase in Fixed Assets + Decrease in NWC
OCF = -2355000 + 740000 - 125000 = -1,740,000

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well